# Parametric Operator Inference to Simulate the Purging Process in Semiconductor Manufacturing

### v1

@@ 1 Introduction
Semiconductors have complex structures that are manufactured through a series of sophisticated processes, such as photolithography for pattern transfer, deposition for thin layer addition, etching for material removal, and doping to modify electrical properties. Across many fabrication steps, particle contamination has long been a major challenge in semiconductor manufacturing. If particles, which typically range in size from a few nanometers to several micrometers, land on the wafer surface, they disrupt the formation of the desired semiconductor structure, significantly reducing both yield and product quality. Figure 1 provides an illustration of the impact of particle contamination on subsequent processes in wafer manufacturing. It shows that particles settling on a patterned surface interfere with critical fabrication steps, such as deposition, where thin films are formed on the substrate, and etching, which selectively removes material to create a desired pattern. Specifically, particle-covered regions prevent material accumulation in the deposition process, while these areas remain unaffected in the etching process. Particle contamination in semiconductor manufacturing can originate from various sources, such as wafer transfer, equipment wear, manufacturing processes, airborne particles in the environment, etc. Among these, the particles generated during the semiconductor manufacturing process itself are the most unavoidable and high-risk sources. This is because the wafer surface must be directly exposed to an environment where various reactions occur to form the desired structures. Therefore, it is important to ensure that particles produced during the manufacturing process do not contaminate the wafer surface and are removed from the chamber.
@@
The plasma-enhanced chemical vapor deposition (PECVD) process is a widely used semiconductor fabrication technique for depositing thin films onto a wafer surface. This process involves using plasma to enhance the chemical reaction of gaseous precursors, which allows the deposition of material at lower temperatures compared to traditional CVD. Especially, particles generated during the PECVD process tend to remain trapped at a certain height above the wafer due to the influence of the electric field while the plasma is active [1]. When the plasma is deactivated, the electric field dissipates, and the purging process—a subprocess of PECVD—removes these particles from the chamber using gas flow. Since this flow primarily drives the movement of particles, accurate analysis of the purging flow field during the purging process is crucial for controlling and preventing particle contamination in the PECVD process.
@@
To accurately identify and address particle contamination mechanisms, both experimental and numerical analyses are essential. Over the years, many studies have leveraged experimental analysis to improve particle contamination control, focusing on contamination mechanisms [2, 3, 4], particle detection during manufacturing [5, 6], and particle removal [7, 8]. Although experimental approaches offer direct and immediate insights into particle contamination control, their feasibility is often limited in real industrial settings due to the highly sensitive operating environment. Numerical simulation-based analysis, on the other hand, allows for a wider range of interrogations of the process and can provide detailed and comprehensive insights into the contamination mechanisms. In particular, computational fluid dynamics (CFD) simulations have been widely used to analyze the fluid flow dynamics within manufacturing equipment, which plays a crucial role in governing particle behavior [9, 10]. However, the substantial computational cost of CFD simulations poses a challenge for their application in semiconductor manufacturing, where rapid problem solving is essential. Recently, an increasing number of studies have developed machine learning models for semiconductor manufacturing by utilizing both experimental and simulated data [11, 12, 13, 14, 15, 16]. These methods either uncover hidden patterns and provide insights for informed decision-making or minimize the turnaround time for analysis to enable faster decision-making. Specifically in the context of particle control, the authors in [17] develop machine learning models trained on experimental data. However, the development of machine learning models trained on numerical simulation data for particle control remains limited.
@@
Our objective is to rapidly and accurately predict the flow field within the PECVD chamber during the purging process, under varying process parameters. This enables effective prediction of particle behavior, which is conducive to minimizing contamination. In this paper, we aim to mitigate the computational bottleneck of conventional CFD simulation by proposing a data-driven reduced-order model (ROM)—a surrogate for a full-order model (FOM)—using data obtained from CFD simulations. In particular, we use Operator Inference (OpInf) [18, 19]—a scientific machine learning method that learns from data a low-dimensional representation of a high-dimensional system—to predict fluid dynamics within the PECVD process chamber. This nonintrusive approach builds a surrogate model without requiring access to FOM numerical operators or routines of the CFD code. More specifically, we build parametric ROMs to forecast the flow field within the chamber by interpolating between models with different purging process parameters. The results demonstrate that parametric OpInf achieves high predictive accuracy for flow field prediction across varying parameters while significantly improving computational efficiency.
@@ 2.1 Computational Domain of the PECVD Chamber
We simulate the purging process in a simplified chamber structure, as illustrated in Figure 2(a). Given that this chamber has a cylindrical shape with a total radius of 200 mm, its physical properties and governing equations remain unchanged across symmetric planes. This allows us to model only one quarter of the chamber, reducing computational cost in the FOM while still capturing the essential flow behavior. We, thus, use symmetric boundary conditions. The top surface, with a radius of 150 mm, serves as the argon (Ar) gas mass flow inlet, while the bottom surface, which covers the radius range of 35 to 65 mm, acts as the pressure outlet. A wafer heater, positioned 8 mm below the top surface, has a radius of 160 mm and supports the wafer during the purging process. We generate a total of 249,262 mesh elements that consist of 232,304 8-node hexahedral elements and 1,448 6-node wedge elements.
Figure 2: (a) Computational domain of a simplified PECVD chamber. B.C.1: mass flow inlet boundary condition, B.C.2: pressure outlet boundary condition, B.C.3: symmetric boundary condition. The side view represents the $xy$-vertical plane where $z=0$ mm. (b) Profiles of the mass flow inlet, pressure outlet, and heater surface temperature over time $t\in[0,1]s$.
@@ 2.2 CFD Simulation of the Purging Process
We simulate the purging process within the PECVD chamber using the commercial software ANSYS Fluent 2023 R1, which solves the governing equations for mass, momentum, and energy using the finite volume method. The flow is assumed to be laminar, and the energy equation is included to capture the effects of heat transfer. When the purging process begins, the Ar mass flow inlet and the pressure outlet boundary conditions change linearly over time, as shown in Figure 2(b). These time-dependent variations are implemented using the boundary condition profile in ANSYS Fluent, while the heater surface is maintained at a constant temperature to ensure stable thermal conditions. To predict the internal flow field within the chamber depending on variations of the inlet mass flow and the outlet pressure, we conduct FOM simulations across a range of two parameter values, denoted as $\mu_{q}$ and $\mu_{p}$. These parameters define the scaling of the final inlet mass flow rate and outlet pressure relative to their initial values. The final inlet Ar mass flow rate (see Figure 2(b)) is given by $Q_{f}=\mu_{q}Q_{0}$, and the final outlet pressure by $P_{f}=\mu_{p}P_{0}$, where $Q_{0}$ and $P_{0}$ are the initial inlet Ar mass flow rate and outlet pressure, respectively. Figure 3 shows the initial flow field in the PECVD chamber. When the purging process begins, steady airflow is formed, with the Ar gas entering through the upper inlet and exiting through the bottom exhaust.
@@ 2.3 Parameter-dependent Full-order Model
Consider a parameter vector $\boldsymbol{\mu}\in\mathcal{M}\subset^{d_{p}}$ of dimension $d_{p}\in\mathbb{N}$. Given the two parameters $\mu_{q}$ and $\mu_{p}$ discussed in Sec. 2.2, our parametric dimension is $d_{p}=2$. The semi-discretization of the nonlinear governing partial differential equations (PDEs) results in an $N$-dimensional system of nonlinear ordinary differential equations (ODEs) written generically as
@@ 2.4 Full-order Model Data from CFD Simulation
The high-fidelity data consists of $n_{x}=249,262$ semi-discretized spatial elements and $n_{v}=5$ variables (pressure, temperature, three velocity components) at each time step and each combination of parameters. Thus, $N=n_{x}\cdot n_{v}=1{,}246{,}310$ becomes the degrees of freedom of the high-fidelity model. The simulation is conducted over $T=1s$, with a constant time step $\delta=0.005s$, resulting in $K=200$ time steps during the purging process. We generate a total of $25$ FOM solutions for five distinct values for both the flow rate parameter $\mu_{q}$ and the pressure parameter $\mu_{p}$. From the FOM data, $d=9$ solutions are used for training, while 16 are used for testing. Given a set of training parameter vectors $\{{\boldsymbol{\mu}_{\ell}}\}_{\ell=1}^{d}$, where $\boldsymbol{\mu}_{\ell}\in\mathcal{M}$ for each $\ell=1,\ldots,d$, we collect the state trajectories for all $d$ parameter instances, and store them in the global data matrix
@@ 3 Parametric Operator Inference for nonintrusive learning of the PECVD purging process
Since we simulate the purging process via ANSYS Fluent, extracting the high-fidelity operators and functions of the semi-discretized PDE is not feasible. For such a setting, Operator Inference (OpInf) has emerged as an efficient scientific machine learning method that constructs predictive polynomial ROMs of high-dimensional dynamical systems [18, 19]. We apply parametric OpInf to interpolate across multiple datasets generated from simulations with varying parameter values, where the parameters $\mu_{q}$ and $\mu_{p}$ represent the scaling factors between the initial and final values of the time-varying boundary conditions. Previous studies have incorporated parametric dependencies into the OpInf framework using different approaches: [20] uses an affine formulation to embed the parametric structure of the governing equations directly into the regression problem, [21] assumes an affine dependence of the reduced operators on the parameters of interest, and [22] utilizes a Taylor series expansion to represent the parametric dependence of the reduced operators. However, since the problem considered herein does not exhibit affine dependence, we adopt an interpolatory regression approach, as we discuss next.
@@ 3.1.2 Parametrized Quadratic ROM
The OpInf framework learns a ROM in polynomial form, and choosing a quadratic polynomial is often a good trade-off between model expressiveness and computational runtime for the ROM simulation [23, 24, 25, 22, 26, 27, 28]. Thus, we opt to learn a parameter-dependent quadratic ROM
$\frac{\mathrm{d}\widehat{\mathbf{s}}(t;\boldsymbol{\mu})}{\mathrm{d}t}$ $=\widehat{\mathbf{c}}(\boldsymbol{\mu})+\widehat{\mathbf{A}}(\boldsymbol{\mu})\widehat{\mathbf{s}}(t;\boldsymbol{\mu})+\widehat{\mathbf{H}}(\boldsymbol{\mu})\left(\widehat{\mathbf{s}}(t;\boldsymbol{\mu})\otimes\widehat{\mathbf{s}}(t;\boldsymbol{\mu})\right)+\widehat{\mathbf{B}}(\boldsymbol{\mu})\mathbf{u}(t;\boldsymbol{\mu}).$
@@ 3.1.3 ROM Learning via Linear Least-squares Regression
$\min_{\widehat{\mathbf{c}}_{\ell},\widehat{\mathbf{A}}_{\ell},\widehat{\mathbf{H}}_{\ell},\widehat{\mathbf{B}}_{\ell}}\sum_{\ell=1}^{d}\sum_{k=0}^{dK-1}\Big{\|}\widehat{\mathbf{c}}_{\ell}+\widehat{\mathbf{A}}_{\ell}\widehat{\mathbf{s}}(t_{k};\boldsymbol{\mu}_{\ell})+\widehat{\mathbf{H}}_{\ell}\big{(}\widehat{\mathbf{s}}(t_{k};\boldsymbol{\mu}_{\ell})\otimes\widehat{\mathbf{s}}(t_{k};\boldsymbol{\mu}_{\ell})\big{)}+\widehat{\mathbf{B}}_{\ell}\mathbf{u}(t_{k};\boldsymbol{\mu}_{\ell})-\dot{\widehat{\mathbf{s}}}(t_{k};\boldsymbol{\mu}_{\ell})\Big{\|}_{2}^{2},$ (3)
where $\widehat{\mathbf{c}}_{\ell}=\widehat{\mathbf{c}}(\boldsymbol{\mu}_{\ell})$, $\widehat{\mathbf{A}}_{\ell}=\widehat{\mathbf{A}}(\boldsymbol{\mu}_{\ell})$, $\widehat{\mathbf{H}}_{\ell}=\widehat{\mathbf{H}}(\boldsymbol{\mu}_{\ell})$, and $\widehat{\mathbf{B}}_{\ell}=\widehat{\mathbf{B}}(\boldsymbol{\mu}_{\ell})$. In the previous equation, $\dot{\widehat{\mathbf{s}}}(t_{k};\boldsymbol{\mu}_{\ell})$ represents the time derivative of the reduced state $\widehat{\mathbf{s}}(t;\boldsymbol{\mu}_{\ell})$ at time $t=t_{k}$, and it can be computed using a suitable time derivative approximation scheme. Equation 3 can be written compactly as
@@
where the reduced operators matrix $\widehat{\mathbf{O}}_{\ell}=\widehat{\mathbf{O}}(\boldsymbol{\mu}_{\ell})=[\widehat{\mathbf{c}}_{\ell}\hskip 5.0pt\widehat{\mathbf{A}}_{\ell}\hskip 5.0pt\widehat{\mathbf{H}}_{\ell}\hskip 5.0pt\widehat{\mathbf{B}}_{\ell}]\in^{r\times d(r,m)}$ and $d(r,m)=1\leavevmode\nobreak\ +\leavevmode\nobreak\ r+r(r+1)/2+m$. The data matrix $\mathbf{D}_{\ell}=[\mathbf{1}_{dK}\hskip 5.0pt\widehat{\mathbf{S}}_{\ell}^{\top}\hskip 5.0pt(\widehat{\mathbf{S}}_{\ell}\otimes\widehat{\mathbf{S}}_{\ell})^{\top}\hskip 5.0pt\mathbf{U}_{\ell}^{\top}]\in^{dK\times d(r,m)}$, where $\mathbf{1}_{dK}\in\mathbb{R}^{dK}$ is a column vector of length $dK$ with all entries set to one, $\widehat{\mathbf{S}}_{\ell}=\widehat{\mathbf{S}}(\boldsymbol{\mu}_{\ell})=[\widehat{\mathbf{s}}(t_{0};\boldsymbol{\mu}_{\ell})\cdots\widehat{\mathbf{s}}(t_{dK-1};\boldsymbol{\mu}_{\ell})]\in^{r\times dK}$, and $\mathbf{U}_{\ell}=\mathbf{U}(\boldsymbol{\mu}_{\ell})=[\mathbf{u}(t_{0};\boldsymbol{\mu}_{\ell})\cdots\mathbf{u}(t_{dK-1};\boldsymbol{\mu}_{\ell})]\in^{m\times dK}$. Also, $\dot{\widehat{\mathbf{S}}}_{\ell}\in^{r\times dK}$ is a matrix whose columns are $\dot{\widehat{\mathbf{s}}}(t_{k};\boldsymbol{\mu}_{\ell})$. We then rewrite the cost function $L(\widehat{\mathbf{O}}_{\ell})$ in (4) as
@@
To minimize $L(\widehat{\mathbf{O}}_{\ell})$, we apply the first-order optimality condition, which involves taking the gradient of $L(\widehat{\mathbf{O}}_{\ell})$ with respect to each $\widehat{\mathbf{O}}_{\ell}$ and setting it to zero. The gradient of each term in the cost function with respect to $\widehat{\mathbf{O}}_{\ell}$ is simplified as
$\nabla_{\widehat{\mathbf{O}}_{\ell}}f(\widehat{\mathbf{O}}_{\ell})$ $=\nabla_{\widehat{\mathbf{O}}_{\ell}}\text{tr}\left(\widehat{\mathbf{O}}_{\ell}\mathbf{D}_{\ell}^{\top}\mathbf{D}_{\ell}\widehat{\mathbf{O}}_{\ell}^{\top}-\widehat{\mathbf{O}}_{\ell}\mathbf{D}_{\ell}^{\top}\dot{\widehat{\mathbf{S}}}_{\ell}^{\top}-\dot{\widehat{\mathbf{S}}}_{\ell}\mathbf{D}_{\ell}\widehat{\mathbf{O}}_{\ell}^{\top}+\dot{\widehat{\mathbf{S}}}_{\ell}\dot{\widehat{\mathbf{S}}}_{\ell}^{\top}\right)$
$=\nabla_{\widehat{\mathbf{O}}_{\ell}}\left[\text{tr}\left(\widehat{\mathbf{O}}_{\ell}\mathbf{D}_{\ell}^{\top}\mathbf{D}_{\ell}\widehat{\mathbf{O}}_{\ell}^{\top}\right)-\text{tr}\left(\widehat{\mathbf{O}}_{\ell}\mathbf{D}_{\ell}^{\top}\dot{\widehat{\mathbf{S}}}_{\ell}^{\top}\right)-\text{tr}\left(\dot{\widehat{\mathbf{S}}}_{\ell}\mathbf{D}_{\ell}\widehat{\mathbf{O}}_{\ell}^{\top}\right)+\text{tr}\left(\dot{\widehat{\mathbf{S}}}_{\ell}\dot{\widehat{\mathbf{S}}}_{\ell}^{\top}\right)\right]$
$=\widehat{\mathbf{O}}_{\ell}\left(\mathbf{D}_{\ell}^{\top}\mathbf{D}_{\ell}+\mathbf{D}_{\ell}^{\top}\mathbf{D}_{\ell}\right)-\left(\mathbf{D}_{\ell}^{\top}\dot{\widehat{\mathbf{S}}}_{\ell}^{\top}\right)^{\top}-\dot{\widehat{\mathbf{S}}}_{\ell}\mathbf{D}_{\ell}+0$
$=2\left(\widehat{\mathbf{O}}_{\ell}\mathbf{D}_{\ell}^{\top}\mathbf{D}_{\ell}-\dot{\widehat{\mathbf{S}}}_{\ell}\mathbf{D}_{\ell}\right).$

### v2

@@ 1 Introduction
have complex structures that are manufactured through a series of sophisticated processes, such as photolithography for pattern transfer, deposition for thin layer addition, etching for material removal, and doping to modify electrical properties. Across many fabrication steps, particle contamination has long been a major challenge in semiconductor manufacturing. If particles, which typically range in size from a few nanometers to several micrometers, land on the wafer surface, they disrupt the formation of the desired semiconductor structure, significantly reducing both yield and product quality. Fig. 1 provides an illustration of the impact of particle contamination on subsequent processes in wafer manufacturing. It shows that particles settling on a patterned surface interfere with critical fabrication steps, such as deposition, where thin films are formed on the substrate, and etching, which selectively removes material to create a desired pattern. Specifically, particle-covered regions prevent material accumulation in the deposition process, while these areas remain unaffected in the etching process. Particle contamination in semiconductor manufacturing can originate from various sources, such as wafer transfer, equipment wear, manufacturing processes, airborne particles in the environment, etc. Among these, the particles generated during the semiconductor manufacturing process itself are the most unavoidable and high-risk sources. This is because the wafer surface must be directly exposed to an environment where various reactions occur to form the desired structures. Therefore, it is important to ensure that particles produced during the manufacturing process do not contaminate the wafer surface and are removed from the chamber.
@@
The plasma-enhanced chemical vapor deposition (PECVD) process is a widely used semiconductor fabrication technique for depositing thin films onto a wafer surface. This process involves using plasma to enhance the chemical reaction of gaseous precursors, which allows the deposition of material at lower temperatures compared to traditional CVD. Especially, particles generated during the PECVD process tend to remain trapped at a certain height above the wafer due to the influence of the electric field while the plasma is active [Particle_Hiroyuki2009]. When the plasma is deactivated, the electric field dissipates, and the purging process—a subprocess of PECVD—removes these particles from the chamber using gas flow. Since this flow primarily drives the movement of particles, accurate analysis of the purging flow field during the purging process is crucial for controlling and preventing particle contamination in the PECVD process.
@@
To accurately identify and address particle contamination mechanisms, both experimental and numerical analyses are essential. Over the years, many studies have leveraged experimental analysis to improve particle contamination control, focusing on contamination mechanisms [Process_Moriya2005, Process_Jeong2024, Process_Munsu2015, ku2024minimization], particle detection during manufacturing [park2018kernel, Detecting_JianGang2021], and particle removal [CLN_Bakhtari2006, CLN_Harald2013]. Although experimental approaches offer direct and immediate insights into particle contamination control, their feasibility is often limited in real industrial settings due to the highly sensitive operating environment. Numerical simulation-based analysis, on the other hand, allows for a wider range of interrogations of the process and can provide detailed and comprehensive insights into the contamination mechanisms. In particular, high-fidelity computational fluid dynamics (CFD) simulations have been widely used to analyze the fluid flow dynamics within manufacturing equipment, which plays a crucial role in governing particle behavior [CFD_Hojun2021, CFD_Sejin2010]. However, the substantial computational cost of CFD simulations poses a challenge for their application in semiconductor manufacturing, where rapid problem solving is essential. Recently, an increasing number of studies have developed machine learning models for semiconductor manufacturing by utilizing both experimental and simulated data [huang2023survey, 10601235, ding2019machine, ding2021machine, jin2024machine, tsunooka2018high]. These methods either uncover hidden patterns and provide insights for informed decision-making or minimize the turnaround time for analysis to enable faster decision-making. Specifically in the context of particle control, the authors in [ML_Feng2023] develop machine learning models trained on experimental data. However, the development of machine learning models trained on numerical simulation data for particle control remains limited.
@@
Our objective is to rapidly and accurately predict the flow field within the PECVD chamber during the purging process, under varying process parameters. This enables effective prediction of particle behavior, which is conducive to minimizing contamination. In this paper, we aim to mitigate the computational bottleneck of conventional CFD simulation by proposing a data-driven reduced-order model (ROM)—a surrogate for a full-order model (FOM)—using data obtained from CFD simulations. In particular, we use Operator Inference (OpInf) [PEHERSTORFER2016196, kramer2024learning]—a scientific machine learning method that learns from data a low-dimensional representation of a high-dimensional system—to predict fluid dynamics within the PECVD process chamber. This nonintrusive approach builds a surrogate model without requiring access to FOM numerical operators or routines of the CFD code. More specifically, we build parametric ROMs to forecast the flow field within the chamber by interpolating between models with different purging process parameters. The results demonstrate that parametric OpInf achieves high predictive accuracy for flow field prediction across varying parameters while significantly improving computational efficiency.
@@ 2.1 Computational Domain of the PECVD Chamber
We simulate the purging process in a simplified chamber structure, as illustrated in Fig. 2(a). Given that this chamber has a cylindrical shape with a total radius of 200 mm, its physical properties and governing equations remain unchanged across symmetric planes. This allows us to model only one quarter of the chamber, reducing computational cost in the FOM while still capturing the essential flow behavior. We, thus, use symmetric boundary conditions. The top surface, with a radius of 150 mm, serves as the argon (Ar) gas mass flow inlet, while the bottom surface, which covers the radius range of 35 to 65 mm, acts as the pressure outlet. A wafer heater, positioned 8 mm below the top surface, has a radius of 160 mm and supports the wafer during the purging process. We generate a total of 249,262 mesh elements that consist of 232,304 8-node hexahedral elements and 1,448 6-node wedge elements.
Figure 2: (a) Computational domain of a simplified PECVD chamber. B.C.1: mass flow inlet boundary condition, B.C.2: pressure outlet boundary condition, B.C.3: symmetric boundary condition. The side view represents the $xy$-vertical plane where $z\penalty 10000\ =\penalty 10000\ 0$ $m$. (b) Profiles of the mass flow inlet, pressure outlet, and heater surface temperature over time $t\in[0,1]s$.
@@ 2.2 CFD Simulation of the Purging Process
The purging process within the PECVD chamber occurs after plasma deactivation and the completion of deposition. During purging, chemical reactions are negligible, and particle transport is assumed to be driven mainly by hydrodynamic drag from the bulk gas flow. The low operating pressures and flow rates typical of purging conditions yield Reynolds numbers well within the laminar regime. The resulting flow field, characterized by pressure, temperature, and velocity distributions, therefore provides the critical insights necessary to inform particle transport in practice. We simulate the purging process within the PECVD chamber using the commercial software ANSYS Fluent 2023 R1, which solves the governing equations for mass, momentum, and energy using the finite volume method. The flow is assumed to be laminar, and the energy equation is included to capture the effects of heat transfer. When the purging process begins, the Ar mass flow inlet and the pressure outlet boundary conditions change linearly over time, as shown in Fig. 2(b). These time-dependent variations are implemented using the boundary condition profile in ANSYS Fluent, while the heater surface is maintained at a constant temperature to ensure stable thermal conditions. To predict the internal flow field within the chamber depending on variations of the inlet mass flow and the outlet pressure, we conduct FOM simulations across a range of two parameter values, denoted as $\mu_{q}$ and $\mu_{p}$. These parameters define the scaling of the final inlet mass flow rate and outlet pressure relative to their initial values. The final inlet Ar mass flow rate (see Fig. 2(b)) is given by $Q_{f}=\mu_{q}Q_{0}$, and the final outlet pressure by $P_{f}=\mu_{p}P_{0}$, where $Q_{0}$ and $P_{0}$ are the initial inlet Ar mass flow rate and outlet pressure, respectively. Fig. 3 shows the initial flow field in the PECVD chamber. When the purging process begins, steady airflow is formed, with the Ar gas entering through the upper inlet and exiting through the bottom exhaust.
@@ 2.3 Parameter-dependent Full-order Model
Consider a parameter vector $\boldsymbol{\mu}\in\mathcal{M}\subset^{d_{p}}$ of dimension $d_{p}\in\mathbb{N}$. Given the two parameters $\mu_{q}$ and $\mu_{p}$ discussed in Section˜2.2, our parametric dimension is $d_{p}=2$. The semi-discretization of the nonlinear governing partial differential equations (PDEs) results in an $N$-dimensional system of nonlinear ordinary differential equations (ODEs) written generically as
@@ 2.4 Full-order Model Data from CFD Simulation
The high-fidelity data consists of $n_{x}=249,262$ semi-discretized spatial elements and $n_{v}=5$ variables (pressure, temperature, three velocity components) at each time step and each combination of parameters. Thus, $N=n_{x}\cdot n_{v}=1{,}246{,}310$ becomes the degrees of freedom of the high-fidelity model. The simulation is conducted over $T=1s$, with a constant time step $\delta=0.005s$, resulting in $K=200$ time steps during the purging process. We generate a total of $25$ FOM solutions for five distinct values for both the flow rate parameter $\mu_{q}$ and the pressure parameter $\mu_{p}$. From the FOM data, $d=9$ solutions are used for training, while 16 are used for testing.
Given a set of training parameter vectors $\{{\boldsymbol{\mu}_{\ell}}\}_{\ell=1}^{d}$, where $\boldsymbol{\mu}_{\ell}\in\mathcal{M}$ for each $\ell=1,\ldots,d$, we collect the state trajectories for all $d$ parameter instances, and store them in the global data matrix
@@ 3 Parametric Operator Inference for nonintrusive learning of the PECVD purging process
Since we simulate the purging process via ANSYS Fluent, extracting the high-fidelity operators and functions of the semi-discretized PDE is not feasible. For such a setting, Operator Inference (OpInf) has emerged as an efficient scientific machine learning method that constructs predictive polynomial ROMs of high-dimensional dynamical systems [PEHERSTORFER2016196, kramer2024learning]. We apply parametric OpInf to interpolate across multiple datasets generated from simulations with varying parameter values, where the parameters $\mu_{q}$ and $\mu_{p}$ represent the scaling factors between the initial and final values of the time-varying boundary conditions. Previous studies have incorporated parametric dependencies into the OpInf framework using different approaches: [mcquarrie2023nonintrusive] uses an affine formulation to embed the parametric structure of the governing equations directly into the regression problem, [yildiz2021learning] assumes an affine dependence of the reduced operators on the parameters of interest, and [farcas2023parametric] utilizes a Taylor series expansion to represent the parametric dependence of the reduced operators. However, since the problem considered herein does not exhibit affine dependence, we adopt an interpolatory regression approach, as we discuss next.
@@ 3.1.2 Parametrized Quadratic ROM
The OpInf framework learns a ROM in polynomial form, and choosing a quadratic polynomial is often a good trade-off between model expressiveness and computational runtime for the ROM simulation [McQuarrie_2021, Swischuk_2020, qian2022reduced, farcas2023parametric, Zastrow2023, 0x003d183f, Rocha2023]. Thus, we opt to learn a parameter-dependent quadratic ROM
$\frac{\mathrm{d}\widehat{\mathbf{s}}(t;\boldsymbol{\mu})}{\mathrm{d}t}$ $=\widehat{\mathbf{c}}(\boldsymbol{\mu})+\widehat{\mathbf{A}}(\boldsymbol{\mu})\widehat{\mathbf{s}}(t;\boldsymbol{\mu})+\widehat{\mathbf{H}}(\boldsymbol{\mu})\left(\widehat{\mathbf{s}}(t;\boldsymbol{\mu})\otimes\widehat{\mathbf{s}}(t;\boldsymbol{\mu})\right)$
$\quad+\widehat{\mathbf{B}}(\boldsymbol{\mu})\mathbf{u}(t;\boldsymbol{\mu}).$
@@ 3.1.3 ROM Learning via Linear Least-squares Regression
$\min_{\widehat{\mathbf{c}}_{\ell},\widehat{\mathbf{A}}_{\ell},\widehat{\mathbf{H}}_{\ell},\widehat{\mathbf{B}}_{\ell}}\sum_{\ell=1}^{d}\sum_{k=0}^{dK-1}\Big\|\widehat{\mathbf{c}}_{\ell}+\widehat{\mathbf{A}}_{\ell}\widehat{\mathbf{s}}(t_{k};\boldsymbol{\mu}_{\ell})+$ (3)
$\widehat{\mathbf{H}}_{\ell}\big(\widehat{\mathbf{s}}(t_{k};\boldsymbol{\mu}_{\ell})\otimes\widehat{\mathbf{s}}(t_{k};\boldsymbol{\mu}_{\ell})\big)+\widehat{\mathbf{B}}_{\ell}\mathbf{u}(t_{k};\boldsymbol{\mu}_{\ell})-\dot{\widehat{\mathbf{s}}}(t_{k};\boldsymbol{\mu}_{\ell})\Big\|_{2}^{2},$
where $\widehat{\mathbf{c}}_{\ell}=\widehat{\mathbf{c}}(\boldsymbol{\mu}_{\ell})$, $\widehat{\mathbf{A}}_{\ell}=\widehat{\mathbf{A}}(\boldsymbol{\mu}_{\ell})$, $\widehat{\mathbf{H}}_{\ell}=\widehat{\mathbf{H}}(\boldsymbol{\mu}_{\ell})$, and $\widehat{\mathbf{B}}_{\ell}=\widehat{\mathbf{B}}(\boldsymbol{\mu}_{\ell})$. In the previous equation, $\dot{\widehat{\mathbf{s}}}(t_{k};\boldsymbol{\mu}_{\ell})$ represents the time derivative of the reduced state $\widehat{\mathbf{s}}(t;\boldsymbol{\mu}_{\ell})$ at time $t=t_{k}$, and it can be computed using a suitable time derivative approximation scheme. Equation˜3 can be written compactly as
@@
[…]
@@
[…]
$\quad+\dot{\widehat{\mathbf{S}}}_{\ell}\dot{\widehat{\mathbf{S}}}_{\ell}^{\top}\big)$
$=\nabla_{\widehat{\mathbf{O}}_{\ell}}\Big[\text{tr}\big(\widehat{\mathbf{O}}_{\ell}\mathbf{D}_{\ell}^{\top}\mathbf{D}_{\ell}\widehat{\mathbf{O}}_{\ell}^{\top}\big)-\text{tr}\big(\widehat{\mathbf{O}}_{\ell}\mathbf{D}_{\ell}^{\top}\dot{\widehat{\mathbf{S}}}_{\ell}^{\top}\big)$
$\quad-\text{tr}\big(\dot{\widehat{\mathbf{S}}}_{\ell}\mathbf{D}_{\ell}\widehat{\mathbf{O}}_{\ell}^{\top}\big)+\text{tr}\big(\dot{\widehat{\mathbf{S}}}_{\ell}\dot{\widehat{\mathbf{S}}}_{\ell}^{\top}\big)\Big]$
$=\widehat{\mathbf{O}}_{\ell}\big(\mathbf{D}_{\ell}^{\top}\mathbf{D}_{\ell}+\mathbf{D}_{\ell}^{\top}\mathbf{D}_{\ell}\big)-\big(\mathbf{D}_{\ell}^{\top}\dot{\widehat{\mathbf{S}}}_{\ell}^{\top}\big)^{\top}-\dot{\widehat{\mathbf{S}}}_{\ell}\mathbf{D}_{\ell}$
$=2\big(\widehat{\mathbf{O}}_{\ell}\mathbf{D}_{\ell}^{\top}\mathbf{D}_{\ell}-\dot{\widehat{\mathbf{S}}}_{\ell}\mathbf{D}_{\ell}\big).$